The solution of the inequality IIxI-1I ≤ 1/2 are ?


1
Expert's answer
2021-10-21T14:36:35-0400

"-1/2\le|x|-1\le1/2"

"1/2\le|x|\le3/2"


"x\ge 1/2" or "x\le -1/2"

"-3/2\le x\le 3/2"


"x\isin [-3/2,-1/2]\lor [1/2,3/2]"